A spinner has four colours: red, yellow, green, and purple. The experimental probability of spinning a red is 0.2, of spinning a yellow is 0.3, and of spinning a green is 0.4. If the spinner was spun 30 times, how many times did it land on purple?
Answer by greenestamps(13209)



P(purple) = 1-(0.2+0.3+0.4) = 1-0.9 = 0.1

If the spinner was spun 30 times, the number of times it landed on purple was 0.1*30 = 3.